OpenStudy (narutoboy14):

what is the slope of the line passes through the points (-1,13) and (4,2)

OpenStudy (narutoboy14):

will give medal

OpenStudy (mathmale):

This all hinges on the concept of "slope" and the straightforward formula used to find the slope of a line thru 2 given points. Please Google "slope of a line" to find that formula.

OpenStudy (narutoboy14):

Let's take a look at the straight line y = ( 2/3 ) x – 4. Its graph looks like this: To find the slope, we will need two points from the line. Pick two x's and solve for each corresponding y: If, say, x = 3, then y = ( 2/3 )(3) – 4 = 2 – 4 = –2.